Circular No. 6385

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996Y IN ANONYMOUS GALAXY J. Mueller reports her discovery of an apparent supernova (red mag ∼

18.5) at α = 11^{h}21^{m}32^{s}.57, δ = +2^{o}53′14′′.0 (equinox 2000.0), which is 7′′ west and 13′′ north of the center of the host galaxy. SN 1996Y was found on a IIIa-F plate exposed by K. M. Rykoski and Mueller with the 1.2-m Oschin Schmidt telescope on Apr. 10 UT in the course of the second Palomar Sky Survey. I. N. Reid reports that no object appears at the position on a Sky Survey IIIa-F plate taken on 1991 Mar. 17. A spectrum obtained on Apr. 19 by S. G. Djorgovski and R. Gal at the 5-m Hale telescope (+ double spectrograph) suggests that this is indeed a supernova.

A. V. Filippenko, University of California at Berkeley, reports that inspection of a CCD spectrum (range 467–750 nm, resolution 0.7 nm) obtained by M. Eracleous and D. C. Leonard (also of Berkeley) on Apr. 21 UT with the 3-m Shane reflector at Lick Observatory shows that the object is a supernova. The spectral type and phase are uncertain pending calibration of the data.

GRO J1744–28 A. B. Giles, NASA Goddard Space Flight Center and Universities

Space Research Association, and the PCA instrument team report: “The PCA experiment on RXTE has now observed the galactic-center transient GRO J1744–28 (cf. IAUC 6291) on 29 occasions. The source behavior continues as reported before (IAUC 6338), with no sign of the large bursts ceasing. On Apr. 19, the nonburst flux was ∼ 225 mCrab (2–60 keV) or ∼ 11 percent of its late-Jan. level. The rate of decline of the nonburst flux has decreased a little, but the flux has followed a smooth decrease given by flux(mCrab) ∼ 2592 − 22D, where D is the day number in 1996. If the present trend were to continue, the nonburst flux would be close to PCA background levels by early May. The decline in the integrated burst flux has continued to lag behind that for the nonburst flux, and a linear projection places it near zero in late May.”

COMET 22P/KOPFF Total visual magnitude estimates (cf. IAUC 6357): Apr. 18.18 UT, 10.1

(J. M. Trigo, Castellon, Spain, 0.18-m reflector); 20.07, 10.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 April 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green